OpenStudy (anonymous):

An Arc of length 30 meters is formed by a central angle A on a circle of radius 15. The measure of A in degrees (2 decimal points) is ... ? I got 28.65 but I'm not completely sure on that.. Want to check

OpenStudy (jhannybean):

 So following this diagram,we can find the measure of our central angle by the formula \[\large s = \frac{\theta}{360}\cdot (2 \pi r)\]\[\large s =\text{ arc length} = 30 \ m\]\[\large r = \text{radius} = 15 \ cm\]\[\large 30 = \frac{\theta}{360} \cdot (2 \cdot \pi \cdot 15)\] Are you able to solve forthis now?